# 3Research Center for Information Technology Innovation (CITI), Academia Sinica (AS), Taipei 11529, Taiwan

### v1

@@ Abstract
This paper presents an agentic artificial intelligence (AI)-driven double deep Q-network (DDQN) scheduling framework for licensed and unlicensed band allocation in New Radio (NR) sidelink (SL) networks. SL must share licensed spectrum with cellular communications (CC) and unlicensed bands with Wi-Fi, posing significant challenges for coexistence. Unlike prior rule-based or threshold-based methods, the proposed agentic scheduler autonomously perceives queueing dynamics, channel conditions, and coexistence states, and adapts its policy to maintain quality-of-service (QoS). Simulation results show that our framework reduces the blocking rate by up to 87.5% compared to threshold-based scheduling under limited licensed bandwidth. These findings demonstrate the potential of Agentic AI to enable stable, QoS-aware, and adaptive scheduling for future NR SL systems.
@@ I Introduction
Reinforcement learning (RL) [16] offers a natural solution by enabling agents to adaptively learn from interaction with stochastic environments. Deep Q-networks (DQN) [17, 18] have been widely applied to wireless scheduling, but are prone to Q-value overestimation. Double DQN (DDQN) [19] addresses this limitation by decoupling action selection from evaluation, achieving more stable convergence. Recent advances in Agentic Artificial Intelligence (AI) [20, 21] emphasize autonomous and context-driven reasoning, where agents integrate multi-dimensional observations such as queue states, channel conditions, and coexistence dynamics to achieve goal-oriented decisions. Leveraging this paradigm, our DDQN-based scheduler embodies agentic decision-making by perceiving heterogeneous contexts, reasoning across layers, and adapting strategies to maintain QoS under constrained spectrum resources.
Motivated by these trends, we propose an Agentic DDQN-based scheduling framework for NR SL that jointly allocates licensed and unlicensed spectrum. Our design integrates queueing delay, channel quality, Wi-Fi coexistence dynamics, and link-switching stability into the state representation and reward. Unlike static heuristics or binary-reward DRL baselines [22, 23], the proposed framework adapts to traffic load and channel variations, achieving up to $87.5\%$ reduction in blocking compared to threshold-based policies and $82.5\%$ compared to random selection under constrained licensed bandwidth. These results highlight the potential of Agentic DDQN-driven learning to deliver QoS-aware, adaptive, and stable scheduling in congested and coexistence-limited NR SL systems.
@@ II Related Work
  • Agentic AI-driven DDQN Scheduling: We propose the first Agentic DDQN-based framework for licensed and unlicensed band allocation in NR SL networks. The scheduler integrates context-driven reasoning by jointly considering queueing delay, residual spectrum availability, channel conditions, and coexistence dynamics, enabling QoS-aware and adaptive scheduling.
  • Unified Mode-Band Decision Formulation: We define a five-action space spanning CC, SL-L, and SL-U modes across licensed and unlicensed bands, allowing the agent to reason holistically across available spectrum options.
  • QoS-Aware Reward Design: The DDQN agent is trained with a composite reward shaped by achievable link capacity, latency, and switching stability, ensuring decisions are aligned with throughput and fairness objectives rather than static heuristics.
  • Coexistence-Aware Access Modeling: We explicitly incorporate LBT and CSMA/CA contention for SL-U, allowing the agent to adapt to Wi-Fi traffic and maintain stable coexistence in contention-heavy environments.
@@ III System Model
Figure 1: Heterogeneous NR SL and Wi-Fi coexistence scenario with licensed (26/28 GHz) and unlicensed (5 GHz) bands, where the gNB allocates CC, SL-L, and SL-U modes.
@@ III-A Channel Model
Large-scale path loss is modeled using the urban microcell (UMi) street canyon model specified in 3GPP TS 38.901 [8]. For distances smaller than the breakpoint distance $d_{BP}$, the LOS path loss is
$PL_{\text{LOS, 1}}=32.4+21\log_{10}(d_{3D})+20\log_{10}(f_{c}),$ (2)
and for larger distances beyond the breakpoint,
$PL_{\text{LOS, 2}}$ $=32.4+40\log_{10}(d_{3D})+20\log_{10}(f_{c})$
$\quad-(9.5\log_{10}(d_{BP}^{2})+(h_{t}-h_{r})^{2}),$ (3)
where $h_{t}$ and $h_{r}$ are the antenna heights of the transmitter and receiver, respectively, and $d_{BP}=\frac{4h_{t}h_{r}f_{c}}{c}$ is the breakpoint distance, $f_{c}$ represents the carrier frequency (Hz), and $c=3\times 10^{8}(m/s)$ as the speed of light. For NLOS, the effective path loss is the maximum of the LOS path loss and an empirically modeled NLOS path loss, which is
$PL_{NLOS}=$ $22.4+35.3\log_{10}(d_{3D})+21.3\log_{10}(f_{c})$
$-0.3(h_{{r}}-1.5).$ (4)
At the receiver, the achievable transmission rate $R$ is calculated using Shannon’s capacity formula $R=B\log_{2}(1+\text{SNR})$, where $B$ is the allocated bandwidth, and $\text{SNR}=\frac{P_{t}|H|^{2}}{N_{0}B}$ is the received signal-to-noise ratio, with $P_{t}$ as the transmission power and $N_{0}$ as the noise variance.
@@ IV-C Proposed Reward Function
where $B$ represents the bandwidth (in Hz) of the selected transmission band, and $\text{SINR}_{t}$ is the received signal-to-interference-plus-noise ratio corresponding to the selected transmission mode and band. This rate-based reward function incentivizes the agent to prioritize high-throughput, low-congestion links, aligning with Agentic AI’s goal-oriented decision-making paradigm. In contrast to prior binary-reward schemes [22, 23], this formulation incorporates link quality and bandwidth utilization directly into the reward, thereby encouraging QoS-aware decisions.
@@ IV-E Objective Function and Learning Formulation
$\vskip-21.68121pt\pi^{*}$ $=\arg\min_{\pi}\limsup_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\mathbbold{1}_{\{B_{t}\}},$ (5a)
@@
V Proposed DDQN-Based Band Allocation
To tackle the joint mode and band selection problem under dynamic wireless environments, we propose a DDQN-based scheme. In the proposed framework, the agent learns an optimal scheduling policy $\pi^{*}$ by approximating the Q-function through interactions with the environment. The blocking penalty is embedded into the reward structure, allowing the agent to balance throughput maximization and packet blocking minimization while implicitly satisfying the QoS constraint.
@@ V-B Training Procedure
$L(\boldsymbol{\theta})=\mathbb{E}\left[\left(y_{t}^{\text{DDQN}}-Q(s_{t},a_{t};\boldsymbol{\theta})\right)^{2}\right],$ (6)
@@ 20         Update target network: $\boldsymbol{\theta}^{-}\leftarrow\boldsymbol{\theta}$;
Algorithm 1 Proposed DDQN-Based Band Allocation
@@ VII Conclusion
In this paper, we have presented an Agentic DDQN-based licensed and unlicensed band allocation scheme for NR SL networks, aiming to minimize blocking under dynamic spectrum availability. Simulation results show a reduction of blockage by up to $87.5\%$ and $82.5\%$ compared to threshold-based and random policies, respectively, under $500$ Mbps licensed bandwidth. These findings highlight the promise of Agentic AI-driven reinforcement learning as a practical scheduling tool for NR SL, enabling efficient and robust coexistence with Wi-Fi in congested environments. The capacity-aware reward design allowed the agent to balance efficiency and fairness through context-driven reasoning. Future work will extend this framework with real-time traffic prediction and adaptive LBT parameter tuning.

### v2

@@ Abstract
In this paper, we present an agentic double deep Q-network (DDQN) scheduler for licensed/unlicensed band allocation in New Radio (NR) sidelink (SL) networks. Beyond conventional reward-seeking reinforcement learning (RL), the agent perceives and reasons over a multi-dimensional context that jointly captures queueing delay, link quality, coexistence intensity, and switching stability. A capacity-aware, quality of service (QoS)-constrained reward aligns the agent with goal-oriented scheduling rather than static thresholding. Under constrained licensed bandwidth, the proposed design reduces blocking by up to 87.5% versus threshold policies while preserving throughput, highlighting the value of context-driven decisions in coexistence-limited NR SL systems.
@@ I Introduction
Reinforcement learning (RL) [16] offers a natural solution by enabling agents to adaptively learn from interaction with stochastic environments. Deep Q-networks (DQN) [17, 18] have been widely applied to wireless scheduling, but are prone to Q-value overestimation. Double DQN (DDQN) [19] addresses this limitation by decoupling action selection from evaluation, achieving more stable convergence. Recent advances in agentic AI highlight that agents differ from conventional RL by extending beyond scalar reward optimization toward autonomous, context-driven, and goal-aligned reasoning [20]. Following this perspective, our DDQN-based scheduler embodies agentic decision-making by perceiving heterogeneous context, reasoning across layers, and adapting strategies to maintain QoS under constrained spectrum resources.
Motivated by these trends, we develop an agentic DDQN framework for NR SL that jointly allocates licensed and unlicensed spectrum. Our design integrates queueing delay, channel quality, Wi-Fi coexistence dynamics, and link-switching stability into the state and reward formulation, departing from static heuristics or binary-reward DRL baselines [21, 22]. Unlike conventional RL schedulers that typically optimize a single reward with simplified states (e.g., queue length or binary success), our agent leverages multi-dimensional perception with a capacity-aware, QoS-constrained objective. This enables context-driven, goal-oriented reasoning across licensed/unlicensed modes and avoids pathological behaviors such as overusing a single band or oscillating between modes.
As demonstrated in our experiments, the proposed scheduler achieves up to $87.5\%$ and $82.5\%$ reduction in blocking relative to threshold and random policies, respectively, under constrained licensed bandwidth, while preserving throughput. These results highlight the potential of agentic DDQN-driven learning, consistent with the RL-oriented definition in [20], to deliver QoS-aware, adaptive, and stable scheduling in coexistence-limited NR SL systems.
@@ II Related Work
  • Agentic DDQN scheduler: A DDQN agent that perceives multi-dimensional context (queueing, residual spectrum, channel quality, coexistence, switching stability) and reasons toward QoS goals, departing from conventional reward-only RL.
  • Unified mode–band action space: A five-action formulation over CC/SL-L/SL-U and 26/28/5 GHz enables holistic decisions across heterogeneous spectrum options.
  • Goal-aligned learning objective: A capacity-aware reward with an explicit long-term throughput constraint steers policies away from myopic reward chasing and toward stable QoS.
  • Coexistence-aware modeling: LBT/CSMA-CA contention is embedded in the environment, letting the agent adapt SL-U access to incumbent Wi-Fi dynamics.
@@ III System Model
Figure 1: Heterogeneous NR SL and Wi-Fi coexistence scenario with licensed (26/28 GHz) and unlicensed (5 GHz) bands, where the gNB allocates $C$, SL-L, and SL-U modes.
@@ III-A Channel Model
Large-scale path loss is modeled using the urban microcell (UMi) street canyon model specified in 3GPP TS 38.901 [8]. For distances smaller than the breakpoint distance $d_{BP}$, the LOS path loss is $PL_{\text{LOS, 1}}=32.4+21\log_{10}(d_{3D})+20\log_{10}(f_{c})$, and for larger distances beyond the breakpoint, $PL_{\text{LOS, 2}}=32.4+40\log_{10}(d_{3D})+20\log_{10}(f_{c})-[9.5\log_{10}(d_{BP}^{2})+(h_{t}-h_{r})^{2}]$, where $h_{t}$ and $h_{r}$ are the antenna heights of the transmitter and receiver, respectively, and $d_{BP}=\frac{4h_{t}h_{r}f_{c}}{c}$ is the breakpoint distance, $f_{c}$ represents the carrier frequency (Hz), and $c=3\times 10^{8}(m/s)$ as the speed of light. For NLOS, the effective path loss is the maximum of the LOS path loss and an empirically modeled NLOS path loss, which is $PL_{NLOS}=22.4+35.3\log_{10}(d_{3D})+21.3\log_{10}(f_{c})-0.3(h_{{r}}-1.5)$. At the receiver, the achievable transmission rate $R$ is calculated using Shannon’s capacity formula $R=B\log_{2}(1+\text{SNR})$, where $B$ is the allocated bandwidth, and $\text{SNR}=\frac{P_{t}|H|^{2}}{N_{0}B}$ is the received signal-to-noise ratio, with $P_{t}$ as the transmission power and $N_{0}$ as the noise variance.
@@ IV Problem Formulation
The proposed framework departs from conventional RL by combining three design elements: (i) rich perception via a multi-dimensional state, (ii) goal alignment through a capacity-aware reward and throughput constraint, and (iii) a holistic action space across licensed and unlicensed modes. Together, these elements instantiate agentic behavior, where the agent perceives heterogeneous context, reasons about QoS trade-offs, and acts over all feasible mode–band options, rather than simply maximizing a scalar reward.
@@ IV-C Proposed Reward Function
where $B$ represents the bandwidth (in Hz) of the selected transmission band, and $\text{SINR}_{t}$ is the received signal-to-interference-plus-noise ratio corresponding to the selected transmission mode and band. This rate-based reward function incentivizes the agent to prioritize high-throughput, low-congestion links, aligning with agentic AI’s goal-oriented decision-making paradigm. In contrast to prior binary-reward schemes [21, 22], this formulation incorporates link quality and bandwidth utilization directly into the reward, thereby encouraging QoS-aware decisions. This capacity-aware reward, together with the long-term throughput constraint, aligns decisions with QoS objectives and discourages degenerate reward-chasing behaviors, thereby instantiating an agentic learning signal.
@@ IV-E Objective Function and Learning Formulation
$\pi^{*}$ $=\arg\min_{\pi}\limsup_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\mathbbold{1}_{\{B_{t}\}},$ (5a)
@@
V Proposed Agentic DDQN Band Allocation
Our agent retains the standard DDQN update rule and network, while its agentic behavior arises from three elements: enriched perception via a multi-dimensional state, a QoS-aligned reward with constraint, and a unified mode–band action space. Building on these elements, the framework learns an optimal scheduling policy $\pi^{*}$ through interactions with the environment, effectively balancing throughput and blocking under dynamic coexistence.
@@ V-B Training Procedure
$L(\boldsymbol{\theta})=\mathbb{E}\left[\left(y_{t}^{\text{DDQN}}-Q(s_{t},a_{t};\boldsymbol{\theta})\right)^{2}\right],\vskip-7.22743pt$ (3)
@@ 20         Update target network: $\boldsymbol{\theta}^{-}\leftarrow\boldsymbol{\theta}$;
Algorithm 1 Proposed Agentic DDQN Band Allocation
@@ VII Conclusion
In this paper, we propose an agentic DDQN-based scheduler for licensed and unlicensed band allocation in NR sidelink networks. Different from vanilla DDQN applications, the design couples rich perception, a goal-aligned reward, and a unified action space, enabling context-driven decisions. Simulations showed up to $87.5\%$ and $82.5\%$ blocking reduction versus threshold-based and random policies at $500$ Mbps licensed bandwidth, while maintaining throughput. These results highlight the promise of agentic AI for robust coexistence, with future work extending to traffic prediction and adaptive LBT tuning.